Currently browsing tag


TEM status

April 13, 2012. After 3 days of work, the TEM is back online.   The system is conditioned to 200kV and aligned at 100kV and 200kV.  I will work on 300kV over the weekend.  I installed a new filament during the process, so the illumination is brighter than it was.   …

FE-SEM auto brightness/contrast control

March 23, 2012.The auto brightness/contrast control is not working properly.  We can manually adjust the B/C while we are waiting for a repair.  Also, there is a manual control that works like the waveform monitor on the 6400 SEM.  I’ll be happy to show you hope to use it.